# Functional differential equations driven by càdlàg rough paths

### v3

@@ 1. Introduction
Most applications of rough path theory to stochastic differential equations (SDEs) crucially rely on the construction of suitable (random) rough paths. To apply the developed theory on RFDEs to Itô SDEs with constant delay, we show that a càdlàg martingale together with its delayed version can be lifted to a random rough path in the spirit of stochastic Itô integration. The key challenge to obtain the “delayed” rough path is that a martingale together with its delayed version is, in general, not a martingale itself, thus preventing the direct use of stochastic Itô integration. For related constructions of random rough paths above fractional Brownian motions we refer to [NNT08, TT12, BMCR14, CE20]. Consequently, one can apply the continuity of the Itô–Lyons map to derive pathwise stability results for stochastic differential equations with constant delay, which plays an important role in many applications, see e.g. [BDNHP18]. In particular, the map $y\mapsto Y$, mapping the initial condition $y$ to the associated solution $Y$ of an SDE with constant delay, is continuous on the space of controlled paths. This resolves an old observation, pointed out by Mohammed [Moh86], about the non-continuity of the flow of stochastic differential equations with delay, for which the initial condition is in fact an initial path.
@@ 2.1. Essentials on rough path theory
For two vector spaces, the space of linear maps from $E_{1}\to E_{2}$ is denoted by $\mathcal{L}(E_{1};E_{2})$; and we write $C^{l}_{b}=C^{l}_{b}(\mathbb{R}^{m};\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ for the space of $l$-times differentiable (in the Fréchet sense) functions $f\colon\mathbb{R}^{m}\to\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k})$ such that $f$ and all its derivatives up to order $l$ are continuous and bounded. We equip this space with the norm
$\|f\|_{C^{l}_{b}}:=\|f\|_{\infty}+\|\mathrm{D}f\|_{\infty}+\cdots+\|\mathrm{D}^{l}f\|_{\infty},$
where $\|\hskip 0.5pt\cdot\hskip 0.5pt\|_{\infty}$ denotes the supremum norm on the corresponding spaces of operators.
@@ 2.2. Existence and uniqueness
Analogously to Step 2, we can obtain a partition $\mathcal{P}=\{0=\tau_{0}<\tau_{1}<\dots<\tau_{N}=T\}$ and, recalling the definition of $\mathcal{B}_{t}$, define the solution $(Y,Y^{\prime})$ on each of the half-intervals $[t_{i},t_{i+1})$ such that
(2.3) $\|Y^{\prime}\|_{p,[t_{i},t_{i+1})}\leq 4C_{F}+\|y^{\prime}\|_{p,[t_{i},t_{i+1})}$
@@ 3.4. RDEs with constant delay
More precisely, we consider the functional $(F,F^{\prime})\colon\mathcal{V}^{p}_{X}([0,r_{1}];\mathbb{R}^{k})\to\mathcal{V}^{p}_{X}([0,r_{1}];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$, for $(\alpha,\alpha^{\prime})\in\mathcal{V}^{p}_{X}([0,r_{1}];\mathbb{R}^{k\ell})$ given by (3.5), and apply Corollary 3.3 (i) to show that there exists a unique solution to the RDE (3.4) on the interval $[0,r_{1})$.
@@ 4.1. Delayed martingales as rough paths
in the spirit of stochastic Itô integration. Recall, for a martingale $(S_{t})_{t\in[0,T]}$, stochastic Itô integration allows to define the integral $\int_{0}^{t}\varphi_{s}\,\mathrm{d}S_{s}$ if $(\varphi_{t})_{t\in[0,T]}$ is a stochastic process with left-continuous sample paths with right-limits which is adapted to the augmented filtration generated by $(S_{t})_{t\in[0,T]}$. For a comprehensive introduction to stochastic integration see, e.g., [Pro05]. In the following, when writing a stochastic integral, like $\int_{0}^{t}\varphi_{s}\,\mathrm{d}S_{s}$, we will always implicitly refer to the augmented filtration generated by $(S_{t})_{t\in[0,T]}$ if not explicitly stated otherwise.
@@ 4.2. SDEs with delay as random RDEs
    Let $(V,V^{\prime})$ be an adapted stochastic process such that $(V(\omega),V^{\prime}(\omega))\in\mathcal{V}^{p}_{X(\omega)}$ for almost every $\omega\in\Omega$. Then, the rough integral exists and coincides $\mathbb{P}$-almost surely with the stochastic Itô integral, that is
@@
    The solution of the SDE (4.3) driven by $X$, and the solution of the RDE (3.4) driven by $\mathbf{X}$, coincide $\mathbb{P}$-almost surely.
@@
$\sum_{(s,t)\in\mathcal{P}}V_{s}X_{s,t}\to\int_{0}^{T}V_{u-}\,\mathrm{d}X_{u},$
in probability as $|\mathcal{P}|\to 0$, see e.g. [Pro05, Chapter II, Theorem 21], therefore the convergence holds $\mathbb{P}$-almost surely, possibly along some subsequence.
@@ Appendix A Local estimates for rough integration
where the implicit multiplicative constant depends only on $p$. For $(V,V^{\prime})=(F(Y),F^{\prime}(Y,Y^{\prime}))$, using Assumption 2.1 (i), we therefore obtain the estimate. ∎

### v4

@@ 1. Introduction
Most applications of rough path theory to stochastic differential equations (SDEs) crucially rely on the construction of suitable (random) rough paths. To apply the developed theory on RFDEs to Itô SDEs with constant delay, we show that a càdlàg martingale together with its delayed version can be lifted to a random rough path in the spirit of stochastic Itô integration. The key challenge to obtain the “delayed” rough path is that a martingale together with its delayed version is, in general, not a martingale itself, thus preventing the direct use of stochastic Itô integration. For related constructions of random rough paths above fractional Brownian motions we refer to [NNT08, TT12, BMCR14, CE20]. Consequently, one can apply the continuity of the Itô–Lyons map to derive pathwise stability results for stochastic differential equations with constant delay, which plays an important role in many applications, see e.g. [BnDNHP18]. In particular, the map $y\mapsto Y$, mapping the initial condition $y$ to the associated solution $Y$ of an SDE with constant delay, is continuous on the space of controlled paths. This resolves an old observation, pointed out by Mohammed [Moh86], about the non-continuity of the flow of stochastic differential equations with delay, for which the initial condition is in fact an initial path.
@@ 2.1. Essentials on rough path theory
For two vector spaces, the space of linear maps from $E_{1}\to E_{2}$ is denoted by $\mathcal{L}(E_{1};E_{2})$; and we write $C^{k}_{b}=C^{k}_{b}(\mathbb{R}^{m};\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{n}))$ for the space of $k$-times differentiable (in the Fréchet sense) functions $f\colon\mathbb{R}^{m}\to\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{n})$ such that $f$ and all its derivatives up to order $k$ are continuous and bounded. We equip this space with the norm
$\|f\|_{C^{k}_{b}}:=\|f\|_{\infty}+\|\mathrm{D}f\|_{\infty}+\cdots+\|\mathrm{D}^{k}f\|_{\infty},$
where $\mathrm{D}^{r}f$ denotes the $r$-th order derivative of $f$, and $\|\hskip 0.5pt\cdot\hskip 0.5pt\|_{\infty}$ denotes the supremum norm on the corresponding spaces of operators.
@@ 2.2. Existence and uniqueness
Analogously to Step 2, we can obtain a partition $\mathcal{P}=\{0=\tau_{0}<\tau_{1}<\dots<\tau_{N}=T\}$ and define the solution $(Y,Y^{\prime})$ on each of the half-intervals $[\tau_{i},\tau_{i+1})$, $i=0,1,\dots,N-1$. We recall the definition of $\mathcal{B}_{t}$ and note that the defining estimates also hold in terms of $y$ since the $p$-variation is invariant under additive shifts, thus, $\|y_{i}\|_{p,[\tau_{i},\tau_{i+1})}=\|y\|_{p,[\tau_{i},\tau_{i+1})}$ and $\|R^{y_{i}}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1})}=\|R^{y}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1})}$. It therefore holds that
(2.3) $\|Y^{\prime}\|_{p,[\tau_{i},\tau_{i+1})}\leq 4C_{F}+\|y^{\prime}\|_{p,[\tau_{i},\tau_{i+1})}$
@@ 3.4. RDEs with constant delay
More precisely, we consider the functional
$(F,F^{\prime})\colon\mathcal{V}^{p}_{X}([0,r_{1}];\mathbb{R}^{k})\to\mathcal{V}^{p}_{X}([0,r_{1}];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k})),$
for $(\alpha,\alpha^{\prime})\in\mathcal{V}^{p}_{X}([0,r_{1}];\mathbb{R}^{k\ell})$ given by (3.5), and apply Corollary 3.3 (i) to show that there exists a unique solution to the RDE (3.4) on the interval $[0,r_{1})$.
@@ 4.1. Delayed martingales as rough paths
in the spirit of stochastic Itô integration. Recall that, for a martingale $(S_{t})_{t\in[0,T]}$ (or, more generally, for a semimartingale $(S_{t})_{t\in[0,T]}$), the stochastic Itô integral $\int_{0}^{t}\varphi_{s}\,\mathrm{d}S_{s}$ can be defined whenever $(\varphi_{t})_{t\in[0,T]}$ is a stochastic process with left-continuous sample paths with right-limits, which is adapted to the augmented filtration generated by $(S_{t})_{t\in[0,T]}$. For a comprehensive introduction to stochastic integration see, e.g., [Pro05]. In the following, when writing a stochastic integral, like $\int_{0}^{t}\varphi_{s}\,\mathrm{d}S_{s}$, we will always implicitly refer to the augmented filtration generated by $(S_{t})_{t\in[0,T]}$ if not explicitly stated otherwise.
@@ 4.2. SDEs with delay as random RDEs
    Let $(V,V^{\prime})$ be an adapted stochastic process such that $V$ takes values in $\mathcal{L}(\mathbb{R}^{e};\mathbb{R}^{k})$ and, identifying $V$ with its extension (by zero) to take values in $\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k})$, we assume that $(V(\omega),V^{\prime}(\omega))\in\mathcal{V}^{p}_{X(\omega)}$ for almost every $\omega\in\Omega$. Then, the rough integral exists and coincides $\mathbb{P}$-almost surely with the stochastic Itô integral, that is
@@
    The solution of the SDE (4.3) with coefficient $f$ and driven by $X$, and the solution of the RDE (3.4) with coefficient $f_{0}$ and driven by $\mathbf{X}$, coincide $\mathbb{P}$-almost surely, if $f_{0}$ be the extension (by zero) of $f$ to a function taking values in $\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k})$.
@@
$\sum_{(s,t)\in\mathcal{P}}V_{s}X_{s,t}\to\int_{0}^{T}V_{u-}\,\mathrm{d}Z_{u},$
in probability as $|\mathcal{P}|\to 0$, noting that $V_{s}X_{s,t}=V_{s}Z_{s,t}$ by definition of $V$, see e.g. [Pro05, Chapter II, Theorem 21], therefore the convergence holds $\mathbb{P}$-almost surely, possibly along some subsequence.
@@ Appendix A Local estimates for rough integration
where the implicit multiplicative constant depends only on $p$. Using Assumption 2.1 (i), we therefore obtain the estimate for $(V,V^{\prime})=(F(Y),F^{\prime}(Y,Y^{\prime}))$. ∎